OpenStudy (anonymous):

help please. How can i writef(x)= 3x^2-6x+1 in the form f(x)=a(x-b)^2+c

OpenStudy (anonymous):

you can use either method: completing the square factoring

OpenStudy (anonymous):

your final answer would be 3(x-1)^2-2

OpenStudy (bradely):

3x^2-6x+1 3(x^2-2x)+1 3(x^2-2x+1-1)+1 3(x^2-2x+1)-3+1 3(x-1)^2-2 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

thanks fpr the help

OpenStudy (zzr0ck3r):

\(ax^2+bx+c=0\iff a(x+\frac{b}{2a})^2=-c+a(\frac{b}{2a})^2\)